OpenStudy (abmon98):

Prove this identity 1/(cosA+sinA) +1/(cosA-sinA)=tan2AcosecA

OpenStudy (anonymous):

take LCM and then simplify...

OpenStudy (abmon98):

CosA+sinA+CosA-sinA/Cos^2A-CosAsinA+CosASinA-Sin^2A 2CosA/Cos^2A-Sin^2A=2tanA/1-tan^2A

OpenStudy (abmon98):

2cosA/Cos2A

OpenStudy (shamim):

plz c my attachment

OpenStudy (abmon98):

Thanks @shamim but i didnt understand the part from 2cosA/Cos2A

OpenStudy (shamim):

i took sinA as both numerator n denominator

OpenStudy (shamim):

we can cancell sinA frm denominator n numerator

OpenStudy (shamim):

i think u know sin2A=2sinA.cosA

OpenStudy (shamim):

u know 1/sinA=cosecA

OpenStudy (abmon98):

Yes i Do Know that 1/sinA=cosecA and that sin2A=2SinACosA so @shamim did you simplify the right side into 2cosA/Cos2A=sin2A/Cos2A*1/Sin

OpenStudy (shamim):

ya i need to get sin2A as numerator

OpenStudy (abmon98):

Multiply the left hand side numerator and denominator by SinA 2CosAsinA/Cos2ASinA

OpenStudy (shamim):

ya

OpenStudy (anonymous):

\[\frac{1}{(cosA+sinA)} +\frac{1}{(cosA-sinA)}=\tan2AcosecA \] \[LHS=\frac{1}{(cosA+sinA)} +\frac{1}{(cosA-sinA)}\] \[LHS=\frac{(cosA-sinA) + (cosA+sinA)}{(cosA+sinA) \times (cosA-sinA)} \] \[\huge \rightarrow LHS=\frac{cosA-sinA + cosA+sinA}{(\cos^2A-\sin^2A) } \] \[\huge \rightarrow LHS=\frac{2cosA}{\cos2A } \] Multiplying Numerator and denominator by sin A, we find: \[\huge \rightarrow LHS=\frac{2cosA \sin A}{\cos2A \sin A} \] \[\huge \rightarrow LHS=\frac{2\sin AcosA }{\cos2A \sin A} \] \[\huge \rightarrow LHS=\frac{\sin2A }{\cos2A \sin A} \] \[\huge \rightarrow LHS=\frac{\sin2A }{\cos2A } \times \frac{1 }{sinA }\] \[\huge \rightarrow LHS=\tan 2A \times cosec A\] \[\huge \rightarrow LHS=\tan 2A cosec A = RHS\] Hence the identity is Proved.

OpenStudy (anonymous):

@Abmon98

OpenStudy (abmon98):

How Can i give a medal to each of you thank you for your Help @dpasingh and @shamim :D

OpenStudy (shamim):

wc

OpenStudy (anonymous):

@Abmon98 please click on the BEST RESPONSE button besude my name

OpenStudy (abmon98):

i have already chose best response i wish i could give a medal to both of you for your help.

OpenStudy (abmon98):

chosen*